A pendulum bob of mass 80 mg and carrying a charge of 2*10^-8 c is at rest in a uniform, horizontal electric field of 20 kV/m. Find the tension in the thread.

Grade:12

3 Answers

Vikas TU
7 years ago
T = mg only as in vertcal direction only mg acts on mass of bob.
qE force acts horizontally thus net force in x – direction is zero.
 = 80 x !0^-3 x 10 N.
Piyush
7 years ago
As the electric field is in horizontal direction and the bob is positively charged it will move in the direction of the field. Lets define the angle made by the bob after displacement from y-axis as alpha from the y axis. Resolving the components we have.Tsin alpha = qE 1Tcos alpha = mg 2Dividing 1by 2 , we have Tanalpha = qE/mg Mass= 80 * 10^-6 kg. g = 9.8m/sq= 2*10^-8. E= 2*10000 Solving the above equation we have Tan alpha = .5012 (approx)Alpha = 27degrees Now from 1T= mg/sinalpha putting the values and solvingT = 8.8*10^-4 newton(approximately) Hope it helps..cheers..
Kushagra Madhukar
3 years ago
Dear student,
Please find the attached solution to your problem.
 
Tension, T = qE/sinθ = 2 x 10-8 x 2 x 104 / sin27o = 8.8 x 10-4 V
 
Hope it helps.
Thanks and regards,
Kushagra